OpenStudy (anonymous):

Factor completely: -6x4y5 - 15x3y2 + 9x2y3

OpenStudy (phi):

These are not too hard if you do the following: the the letters raised to a power, the letter has to be in every term. If it is, take the one with the smallest power. For example, in this one you have x^4 x^3 and x^2 so x is in every term. choose x^2 as part of your GCF. for the numbers, break them into their primes e.g. 6= 2*3 15= 3*5 9=3*3 look for anything in common (you can treat the numbers like letters if you want and use the rule up above)

OpenStudy (anonymous):

\[-3 x^2 y^2 \left(2 x^2 y^3+5 x-3 y\right) \]